Craig is 24 years younger than his mother. Together their ages total 56 years. How is each person? Explain how you found your answer. I need help why this question i have been trying for one hour to do it.
Found 3 solutions by John10, Gogonati, lwsshak3:
Answer by John10(297)

Let x be Craig's age and y be his mom's age
The first sentence tell us that:
y = x + 24
Their sum of ages is 56:
x + y = 56
You need to solve the system of two linear equations:
y = x + 24 (1)
x + y = 56 (2)
Substitute (1) into (2):
x + (x + 24) = 56
2x + 24 = 56
2x = 32
x = 32/2
x = 16
then you find y :
x + y = 56
16 + y = 56
y = 56 - 16
y = 40
Thus Craig's age is 16 yrs old
His mom is 40 yrs old
John10:)

Answer by Gogonati(855)

Let mother's age be x years , then Craig's age is x-24 years. Since together their ages total is 56 years, we write the equation:
x+x-24=56, solving this equation we get mother's age.
2x-24=56
2x-24+24=56+24
2x=80 divide both sides by 2, and get x=40.
Answer:Mother's age is 40, and Craig's age is 40-24= 16.

Answer by lwsshak3(11628)

Craig is 24 years younger than his mother. Together their ages total 56 years. How is each person? Explain how you found your answer
***
let x= mother's age
x-24=Craig's age
ages of mother and son total 56
x+(x-24)=56
x+x-24=56
2x=80
x=40
ans:
mother's age=40
Craig's age=40-24=16